OpenStudy (anonymous):

Ultraviolent rays, which cause sun burn by penetrating only the top level of kin, have a wavelength about 1000 times the wavelength of an X-ray. Find the wavelength of ultraviolent rays. Show your work.

OpenStudy (anonymous):

Medical X-rays, with a wavelength of about 10 ^-10m, can penetrate completely through you skin.

OpenStudy (anonymous):

can u help me...

OpenStudy (frostbite):

Why don't you just multiply the wavelength of an X-ray with 1000 in order to get the wavewength of a UV ray?

OpenStudy (frostbite):

Not sure, but sounds like that it is only that.

OpenStudy (frostbite):

beside we know that X-rays have a wavelength on about 100 nano meter.. if we take 10^-10 meter and convert to nano meters.. we get 0.1 nm. Multiply that with 1000 and you get 100 nm.

OpenStudy (anonymous):

ok so 10^-10m * 1000

OpenStudy (frostbite):

if I had to guess, yes. Because we know the multiplication factor and one of the wavelengths. However if I were you I would calculate in nano meters. nano = 10^-9

OpenStudy (anonymous):

ok this wat i got = 1000000000000

OpenStudy (frostbite):

wait.. how did you get that huge number?

OpenStudy (anonymous):

i got t bye mulitplyn 10^-9*1000

OpenStudy (frostbite):

I doubt that. Lets do it in the head: \[10^{-9}*1000=10^{-9}*10^{3}=10^{-9+3}=10^{-6}\]

OpenStudy (anonymous):

oh ok so the answer is 10^-6

OpenStudy (frostbite):

wait...

OpenStudy (frostbite):

it was not 10^-9 but 10^-10

OpenStudy (frostbite):

\[10^{-10} m*1000=10^{-10} m *10^{3}=10^{-10+3} m=10^{-7} m\] \[10^{-7} m * 10^{9} \frac{ nm }{ m }=100 nm\]

OpenStudy (anonymous):

so 100nm is that right

OpenStudy (frostbite):

Yes 100 nano meters :)

OpenStudy (frostbite):

I think. not 100% sure if the procedure is right... seems to esay for some reason.

OpenStudy (anonymous):

ok so 100nm, can you help me with another one

OpenStudy (frostbite):

Just post it then I might take a look :)

OpenStudy (frostbite):

I ment as a question :)

OpenStudy (anonymous):

Consider the sequence 2, 6, 18, 54…. Let n= the term number in the sequence. Let A (n) = the value of the nth term of the sequence. 4. What is the common ratio of the sequence? Common ratio: 5. Complete each statement. a. A (1) =2=2*3^? b. A (2) =6=2*3=2*3^? c. A (3) =18=2*3*3=2*3^? d. A (4) =54=2*3*3*3=2*3^? 6. What is the relationship between the exponent of the base 3 and the value of n? 7. Complete the statement: A (n) = 2*3^?

OpenStudy (frostbite):

Hmmm I think you need to post this as a seperate question, not sure I fully understand the question.

OpenStudy (anonymous):

Activity 2 Consider the sequence 2, 6, 18, 54…. Let n= the term number in the sequence. Let A (n) = the value of the nth term of the sequence. 4. What is the common ratio of the sequence? Common ratio: 5. Complete each statement. a. A (1) =2=2*3^? b. A (2) =6=2*3=2*3^? c. A (3) =18=2*3*3=2*3^? d. A (4) =54=2*3*3*3=2*3^?

OpenStudy (frostbite):

I just ment i don't understand the way they put it... I'm not from the US so I have to translate the words back and forth, and when I do so I get something that don't make sence. All I can say is that you can discribe your function/seq as this: A(n)=2*3^n, n belongs to all whole positive numbers.

OpenStudy (anonymous):

ok, is that is that all you get. @Frostbite wat bout you @aspenfields

OpenStudy (anonymous):

im just thinking and trying to figure it out atm

OpenStudy (anonymous):

haye ca you put that work back to the first one please...@Frostbite

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

ok while we are waitin for that one can u help me with this other one...please!!!

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

A. What is the common ratio of the sequence 80, 20, 5, 5/4…? Common ratio: B. What are the next three terms in the sequence? Next three terms:

OpenStudy (anonymous):

haye can you please help.... @skullpatrol

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

ok are you going tew help me on the first one or the next three term, but with the first one o idea and the second one i do i have 1 idea

OpenStudy (skullpatrol):

What number do you need to multiply by 80 to get 20?

OpenStudy (anonymous):

you cant muliply anything by 80 to get 20 but 20*4 is 80

OpenStudy (skullpatrol):

How about $$\frac{1}{4}$$

OpenStudy (anonymous):

oh so 80 tmes 1/4 is 20

OpenStudy (skullpatrol):

$$80*\frac{1}{4}=20$$

OpenStudy (anonymous):

oh ok so 20*1/4=5

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

thanks so much ok so can you help me on the other one now.... the first one or do you want me to post it up there

OpenStudy (anonymous):

wait can you help me on the next three terms

OpenStudy (skullpatrol):

Just keep multiplying by 1/4...

OpenStudy (anonymous):

ok is the next term 5/16, then 5/64, then 5/256

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

are yall still there.@skullpatrol @tayluvsya

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

so i did it right. ok so can you help me on the first one

OpenStudy (anonymous):

i really dont get so can you please help me

OpenStudy (skullpatrol):

What was the question?

OpenStudy (anonymous):

Activity 2 Consider the sequence 2, 6, 18, 54…. Let n= the term number in the sequence. Let A (n) = the value of the nth term of the sequence. 4. What is the common ratio of the sequence? Common ratio: 5. Complete each statement. a. A (1) =2=2*3^? b. A (2) =6=2*3=2*3^? c. A (3) =18=2*3*3=2*3^? d. A (4) =54=2*3*3*3=2*3^?

OpenStudy (skullpatrol):

What do you have to multiply 2 by to get 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so all of them are three is that alright

OpenStudy (skullpatrol):

Just like multiplying by 1/4 each time.

OpenStudy (anonymous):

ok that i dont get can you please help me i have to turn this in before 12:30

OpenStudy (skullpatrol):

To get the next term multiply by 3

OpenStudy (anonymous):

so a. 6 b.18 c.54 d. 162 is that right

OpenStudy (skullpatrol):

Yes.

OpenStudy (skullpatrol):

The whole idea behind these type of questions is to answer: "what do I need to multiply by to get the next term"? and then continue using that number.

OpenStudy (anonymous):

so what i got up there are those the right answers....

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

ok i have 3 more

OpenStudy (anonymous):

actually it comes along with it, look back up with the one we just worked out and it says the common ratio, is the ratio times 1/4 or what

OpenStudy (anonymous):

haye wat happened @skullpatrol